a = [ 6  5]
    [-5 -5] and 

b = [ 5   0] 
    [-5  -1]

a. a + b
b. 5a
c. 3a - 2b
d. ab


Found 2 solutions by Edwin McCravy, ikleyn:
Answer by Edwin McCravy(20054)

 a = [ 6  5]
     [-5 -5] 

and 

b = [ 5   0]
    [-5  -1]

a. a + b

[ 6  5]     [ 5   0]    [ 6+5   5+0]   [ 11   5]
[-5 -5]  +  [-5  -1]  = [-5-5  -5-1] = [-10  -6]  



b. 5a

  [ 6  5]   [ 5×6   5×5]   [ 30   25]
5×[-5 -5] = [5×-5  5×-5] = [-25  -25] 

c. 
         
3a - 2b = 

  [ 6  5]       [ 5   0]    [ 3×6   3×5]   [ 2×5   2×0]    
3×[-5 -5]  -  2×[-5  -1]  = [3×-5  3×-5] - [2×-5  2×-1] =  

 
[ 18  15]   [ 10  0]   [    18-10      15-0]   [ 18-10   15-0]  
[-15 -15] - [-10 -2] = [-15-(-10)  -15-(-2)] = [-15+10  -15+2] =

[ 8  15]
[-5 -13]

d. ab

[ 6  5] [ 5   0]    [     6×5+5×(-5)       6×0+5×(-1)]   [30-25  0-5]
[-5 -5]×[-5  -1]  = [(-5)×5+(-5)(-5)  (-5)×0+(-5)(-1)] = [-25+25 0+5] =

[5  -5]  
[0   5]

Edwin



Answer by ikleyn(52777)

.

To add matrices, add their elements in each cell.  

    Do it cell by cell, for each and every cell.




To subtract matrices, subtract their elements in each cell.  

    Do it cell by cell, for each and every cell.




To multiply a matrix by a number, multiply each element of the matrix in its cell by the number.

    Do it cell by cell, for each and every cell.


These simple rules will help you to survive in the matrix world.